# Algebraic depletion interactions in two-temperature mixtures

### v1

@@ Abstract
The phase separation that occurs in two-temperature mixtures, which are driven out of equilibrium at the local scale, has been thoroughly characterized, but much less is known about the depletion interactions that drive it Using numerical simulations in dimension 2, we show that the depletion interactions extend beyond two particle diameters in dilute systems, as expected at equilibrium, and decay algebraically with an exponent $-4$. Solving for the $N$-particle distribution function in the stationary state, perturbatively in the interaction potential, we show that the algebraic correlations arise from triplets of particles at different temperatures, and recover the exponent $-4$.
@@
At equilibrium, coupling two objects to a critical field induces an algebraic effective interaction between these objects. The field may be the electromagnetic field, as in the Casimir effect [1, 2], or the density fluctuations in a binary mixture close to its critical point [3, 4, 5]. In the latter case, inducing such interactions requires the fine tuning of the temperature [5]. Out of equilibrium, algebraic correlations and the effective interactions that they induce are much more frequent and usually do not require fine tuning. They arise for instance in driven systems, such as an electrolyte or a binary colloidal mixture under an electric field [6, 7, 8, 9, 10]. Transient algebraic interactions can also arise during the relaxation following a temperature quench [11, 12]. Finally, active systems can induce interactions [13], which may decay algebraically for asymmetric bodies [14, 15]. Here, we show that algebraic interactions spontaneously appear in mixtures of particles connected to different thermostats, an out of equilibrium system known to exhibit phase separation [16, 17, 18]. This algebraic depletion interaction thus does not require self-propulsion and act between spherical bodies.
@@
Two temperature mixtures are now a paradigmatic model to study non-equilibrium phenomena. They are driven out of equilibrium locally, such as active systems, but do not involve self-propulsion; they are thus simpler that the active-passive mixture that they may represent [19, 20]. Two temperature mixtures can also be mapped to systems with non-reciprocal interactions [21, 22], which have gained attention recently [23]. The hot and cold particles may phase separate even when the interactions between the particles are purely repulsive and identical, provided that the temperature ratio is large enough [17, 18, 24], resulting in dense droplets of cold particles coexisting with a dilute gas of hot particles. Grosberg and Joanny provided an analytical prediction for the instability threshold in a dilute system [17], showing that denser systems are easier to phase separate. Recently, McCarthy et al found that the two species may mix again at high density [25].
@@
To understand the phase separation mechanism, it is instructive to consider the depletion interaction between the cold particles that is induced by the hot particles, which is encoded in the pair correlation of the cold particles [18]. In the approach of Grosberg and Joanny, the free energy of the system can be minimized with respect to the density of the hot particles to get an effective free energy functional for the cold particles only [17]. However, this free energy contains only a coarse-grained description of the interaction between the particles, namely the integral of the Mayer function, and does not provide the shape of the effective interaction. An explicit expression of the effective interaction has been obtained in Ref. [24] using the potential of mean force; however, the application of this approach to an out-of-equilibrium system is questionable. Indeed, we will see that generalizing different equilibrium methods to compute the effective interactions in this out-of-equilibrium system lead to different results, similarly to the situation observed with the pressure in a gas of active particles [26].
@@
In this Letter, we compute the depletion interaction between cold particles that is due to the hot particles analytically, in the dilute limit and perturbatively in the interaction strength $\epsilon$. Restricting ourselves to the dilute limit amounts to consider a system of two cold particles and a single hot one. Contrary to two-body problems [17, 21], this three-body system cannot be mapped onto an effective equilibrium system [27]. At order $\epsilon^{2}$, we show that the potential of mean force argument should be corrected by a purely non-equilibrium three-body potential. At order $\epsilon^{3}$, we find that the three-body potential gives rise to algebraic interactions decaying as $r^{-4}$ in spatial dimension $d=2$. Our results are supported by numerical simulations.
@@
We consider harmonic spheres with positions $\bm{x}_{i}(t)$, connected to thermostats with temperature $T_{i}$ and obeying an overdamped Langevin dynamics:
@@
where $V(r)=\epsilon(1-r)^{2}\theta(1-r)$, $\theta$ being the Heaviside function. The vectorial Gaussian white noises $\bm{\eta}_{i}$ have a unit variance and are uncorrelated. The particles diameter and mobility and the Boltzmann constant have been set to unity. We consider only two species $\alpha\in\{A,B\}$ connected to thermostats $T_{\alpha}$. To focus on the interactions induced by the particles $B$ on the particles $A$, we assume that there are only two $A$ particles $i\in\{1,2\}$ and a small density $\rho$ of $B$ particles. We expand the pair distribution function of the two $A$ particles $g(r)$ as $g(r)=\exp(-V(r)/T_{A})\left[1+\sum_{n=1}^{\infty}\rho^{n}g_{n}(r)\right]$ [28]. We focus on the first term of the expansion, $g_{1}(r)$; for simplicity, we denote it $\hat{g}(r)$ and refer to it as the “correlation”:
@@
Both the full expression (3) and its weak interaction limit (4) are zero beyond two particle diameters, here corresponding to $r=2$. The same expressions can be recovered using the potential of mean force [28] (App. B.1.2).
@@
Alternatively, assuming that the $N$-particle distribution function is given by $f(\bm{X})=\exp\left(-\sum_{\langle ij\rangle}v_{ij}(\bm{x}_{i}-\bm{x}_{j})\right)$, where $\bm{X}=(\bm{x}_{i})_{1\leq i\leq N}$, we find (App. B.2.1)
@@
We simulated Eq. (1) numerically [29] (App. A). In order to improve the statistics, we used the same small density $\rho$ of $A$ and $B$ particles. In this situation, an additional depletion interaction is induced by the particles $A$ themselves, so that $\hat{g}(r)=\hat{g}^{A}(r)+\hat{g}^{B}(r)$, where $\hat{g}^{A}(r)=\hat{g}_{\mathrm{eq}}(r)$ (Eq. (3)). The correlations $\hat{g}(r)$ calculated from simulations with $\epsilon=10$, $\rho=0.05$, $T_{A}=1$ and different values of $T_{B}$ are presented in Fig. 1(a). As expected, we observe an increase of the correlation at contact $\hat{g}(1)$ due to the depletion interaction. This increase scales as $\epsilon^{2}$ with the interaction strength $\epsilon$, as expected from the different theoretical predictions (Fig. 1(b)). However, in contradiction with the predictions, the depletion interaction extends beyond two diameters as soon as the system departs from equilibrium ($T_{A}=T_{B}$). Beyond two particle diameters, the radial dependence of the depletion interaction is compatible with an algebraic decay, $\hat{g}(r)\sim G/r^{4}$. The prefactor $G$ of the algebraic decay increases with the interaction strength as $\epsilon^{3}$ for small $\epsilon$ and saturates for large values of $\epsilon$, corresponding to the hard sphere limit (Fig. 1(b)). The effect of the temperature $T_{B}$ is non-monotonic: $G$ first increases and then decreases at large temperatures (Fig. 1(c)). The temperature where the effect is maximal increases with the interaction strength $\epsilon$. We now return to theory to (i) elucidate the discrepancy between the two arguments adapted out of equilibrium and (ii) find the origin and the characteristics of the algebraic decay.
@@
$\hat{g}(r)=\frac{V*V(r)}{T_{A}T_{AB}}.$ (9)
This expression is in quantitative agreement with the numerical simulations for $r\lesssim 2$ (Fig. 2 and 1(b)). It takes a similar form as the results obtained by adapting equilibrium arguments (Eqs. (6, 8)), but the temperatures enter differently in the prefactor. As the expression (9) is exact at the order $\epsilon^{2}$, we conclude that none of the two equilibrium calculations can be adapted out of equilibrium.
Figure 2: Correlation $\hat{g}(r)$ for $T_{A}=1$, $T_{B}=3$ and $\epsilon=3$: numerical simulations (circles) and theoretical prediction (black solid line) with the contributions of order $\epsilon^{2}$ and $\epsilon^{3}$ (red and blue dashed lines, respectively).
Yet, the expression (9) vanishes at two particle diameters and cannot explain the algebraic decay observed in the simulations (Fig. 1(a)). This is not surprising as it is correct at order $\epsilon^{2}$ while the simulations indicate an $\epsilon^{3}$ dependence of the algebraic decay. We now turn to a small density expansion that is valid for any interaction strength. The $N$-particle distribution $f(\bm{X})$ is stationnary under the Smoluchowski equation describing the microscopic dynamics (1). To isolate the many-body effects from the pair correlation obtained by Grosberg and Joanny [17], we write it as
@@
The depletion interaction encoded in $\hat{g}(r)$ can be computed from the three-body distribution $f_{3}(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3})$. Expanding $f_{3}$ in powers of $\epsilon$ and integrating over $\bm{x}_{3}$, we find (App. D.3.1)
$\hat{g}=\int\text{d}\bm{x}_{3}\left[v_{13}v_{23}-\phi^{(2)}\right.\\ \left.-\frac{1}{2}\left(v_{13}^{2}v_{23}+v_{13}v_{23}^{2}\right)+(v_{13}+v_{23})\phi^{(2)}-\phi^{(3)}\right]\\ +\mathcal{O}(\epsilon^{4}).$ (12)
@@
To compute $\phi^{(3)}$, we plug $\phi^{(2)}$ into the second term of Eq. (11), and balance it with the Laplacian in the first term (App. D.2.2). Using $\phi^{(2)}$ and $\phi^{(3)}$ in the correction (12) leads the depletion interaction at order $\epsilon^{3}$ (App. D.3.3). The resulting expression contains many terms involving double integrals in Fourier space, which can be performed numerically, yielding a good agreement with the $\hat{g}(r)$ obtained from simulations (Fig. 2). In particular, the algebraic decay is recovered. The small wavevector behavior in Fourier space points to an algebraic decay with exponent $-2d$ in spatial dimension $d$ (App. D.4.1). The prefactor can be obtained numerically and is in quantitative agreement with the simulations for weak interactions (Fig. 1(b)). The non-monotonic dependence of the prefactor on $T_{B}$ is also recovered; more precisely, the theory predicts an increase as $(T_{B}-T_{A})$ followed by a decrease as $T_{B}^{-2}$, which are compatible with the results from the simulations (Fig. 1(c)). In contrast, the order $\epsilon^{2}$ decays monotonously with $T_{B}$ (Eq. (9)).
We have shown that a species $B$ at temperature $T_{B}$ induces a depletion interaction decaying as $r^{-4}$ between particles of a species $A$ at temperature $T_{A}\neq T_{B}$ in the dilute limit. We have found that the correction originates from a three-body contribution $\phi$ (Eq. (11)) to the $N$-particle distribution that arises from triplets of particles that are not at the same temperature. Our calculation of $\phi$ is pertubative, but the algebraic decay is observed from weak to strong interactions in our simulations. This suggests that pushing the perturbative computation to higher orders would modify the numerical prefactor of the algebraic decay, but not the exponent.
@@
The depletion interaction that we have unveiled is the driving force behind the formation of dense droplets of cold particles when they separate from the hot particles. As the algebraic tail of the interaction originates from temperature differences ($\tau_{ijk}$ in Eq. (11)), this effect is reversed if one considers the depletion interaction induced by the cold particles on the hot ones, so that the algebraic tail is repulsive in this case. This repulsion between hot particles may work with the attraction between cold particles to trigger the phase separation.

### v2

@@ Abstract
The phase separation that occurs in two-temperature mixtures, which are driven out of equilibrium at the local scale, has been thoroughly characterized, but much less is known about the depletion interactions that drive it. Using numerical simulations in dimension 2, we show that the depletion interactions extend beyond two particle diameters in dilute systems, as expected at equilibrium, and decay algebraically with an exponent $-4$. Solving for the $N$-particle distribution function in the stationary state, perturbatively in the interaction potential, we show that algebraic correlations with an exponent $-2d$ arise from triplets of particles at different temperatures in spatial dimension $d$. Finally, simulations allow us to extend our results beyond the perturbative limit.
@@
At equilibrium, coupling two objects to a critical field induces an algebraic effective interaction between these objects. The field may be the electromagnetic field, as in the Casimir effect [1, 2], or the density fluctuations in a binary mixture close to its critical point [3, 4, 5]. In the latter case, inducing such interactions requires the fine tuning of the temperature [5]. Out of equilibrium, algebraic correlations and the effective interactions that they induce are much more frequent and usually do not require fine tuning. They arise for instance in driven systems, such as an electrolyte or a binary colloidal mixture under an electric field [6, 7, 8, 9, 10]. Transient algebraic interactions can also arise during the relaxation following a temperature quench [11, 12]. Finally, active systems can induce interactions [13], which may decay algebraically for asymmetric bodies [14, 15] or if the active system is flocking [16]. Here, we show that algebraic interactions spontaneously appear in mixtures of particles connected to different thermostats, an out of equilibrium system known to exhibit phase separation [17, 18, 19]. This algebraic depletion interaction thus does not require self-propulsion and acts between spherical bodies.
@@
Two-temperature mixtures are now a paradigmatic model to study non-equilibrium phenomena. They are driven out of equilibrium locally, such as active systems, but do not involve self-propulsion; they are thus simpler than the active-passive mixture that they may represent [20, 21, 22]. Two-temperature mixtures can also be mapped to systems with non-reciprocal interactions [23, 24], which have gained attention recently [25]. The hot and cold particles may phase separate even when the interactions between the particles are purely repulsive and identical, provided that the temperature ratio is large enough [18, 19, 26], resulting in dense droplets of cold particles coexisting with a dilute gas of hot particles. Grosberg and Joanny provided an analytical prediction for the instability threshold in a dilute system [18], showing that denser systems are easier to phase separate. Recently, McCarthy et al found that the two species may mix again at high density [27].
@@
To understand the phase separation mechanism, it is instructive to consider the depletion interaction between the cold particles that is induced by the hot particles, which is encoded in the pair correlation of the cold particles [19]. In the approach of Grosberg and Joanny, the free energy of the system can be minimized with respect to the density of the hot particles to get an effective free energy functional for the cold particles only [18]. However, this free energy contains only a coarse-grained description of the interaction between the particles, namely the integral of the Mayer function, and does not provide the shape of the effective interaction. An explicit expression of the effective interaction has been obtained in Ref. [26] using the potential of mean force; however, the application of this approach to an out-of-equilibrium system is questionable. Indeed, we will see that generalizing different equilibrium methods to compute the effective interactions in this out-of-equilibrium system leads to different results, similarly to the situation observed with the pressure in a gas of active particles [28].
@@
In this Letter, we investigate the depletion interaction between cold particles that is due to the hot particles in the dilute limit using analytical calculations and numerical simulations. Restricting ourselves to the dilute limit amounts to consider a system of two cold particles and a single hot one. Contrary to two-body problems [18, 23], this three-body system cannot be mapped onto an effective equilibrium system [29]. We compute the depletion interaction analytically, perturbatively in the interaction strength $\epsilon$. At order $\epsilon^{2}$, we show that the potential of mean force argument should be corrected by a purely non-equilibrium three-body potential. At order $\epsilon^{3}$, we find that the three-body potential gives rise to algebraic interactions decaying as $r^{-2d}$ in spatial dimension $d$. Simulations confirm our analytical results for weak interactions and show that the algebraic decay is still present for strong interactions, allowing us to discuss the dependence of the magnitude of this decay on the interaction potential.
@@
We consider particles with positions $\bm{x}_{i}(t)$, connected to thermostats with temperature $T_{i}$ and obeying an overdamped Langevin dynamics:
@@
where $V(r)=\epsilon U(r)$ is the interaction potential. The vectorial Gaussian white noises $\bm{\eta}_{i}$ have a unit variance and are uncorrelated. The particles mobility and the Boltzmann constant have been set to unity. We consider only two species $\alpha\in\{A,B\}$ connected to thermostats $T_{\alpha}$. To focus on the interactions induced by the particles $B$ on the particles $A$, we assume that there are only two $A$ particles $i\in\{1,2\}$ and a small density $\rho$ of $B$ particles. We expand the pair distribution function of the two $A$ particles $g(r)$ as $g(r)=\exp(-V(r)/T_{A})\left[1+\sum_{n=1}^{\infty}\rho^{n}g_{n}(r)\right]$ [30]. We focus on the first term of the expansion, $g_{1}(r)$; we denote it $\hat{g}(r)$ and refer to it as the “correlation”:
@@
This order is not affected by the interactions between the $B$ particles, which can render the depletion interaction nonmonotonic [31].
@@
If the potential $U(r)$ has a finite range $\sigma$, which represents the diameter of the particles, both the full expression (3) and its weak interaction limit (4) are zero beyond two particle diameters. The same expressions can be recovered using the potential of mean force [30] (App. B.1.2).
@@
Alternatively, assuming that the $N$-particle distribution function is given by $f(\bm{X})=\exp\left(-\sum_{\langle ij\rangle}v_{ij}(\bm{x}_{i}-\bm{x}_{j})\right)$, where $\bm{X}=(\bm{x}_{i})_{1\leq i\leq N}$, we get (App. B.2.1)
@@
We simulated Eq. (1) numerically [32] (App. A) in spatial dimension $d=2$ for three different interactions between particles with diameter $\sigma=1$: harmonic, $U_{\mathrm{harm}}(r)=(1-r)^{2}\theta(1-r)$, $\theta$ being the Heaviside function, Gaussian, $U_{\mathrm{Gauss}}(r)=\exp(-6r^{2})$, and WCA, $U_{\mathrm{WCA}}(r)=(r^{-6}-1)^{2}\theta(1-r)$. In order to improve the statistics, we used the same small density $\rho$ of $A$ and $B$ particles. In this situation, an additional depletion interaction is induced by the particles $A$ themselves, so that $\hat{g}(r)=\hat{g}^{A}(r)+\hat{g}^{B}(r)$, where $\hat{g}^{A}(r)=\hat{g}_{\mathrm{eq}}(r)$ (Eq. (3)). The correlations $\hat{g}(r)$ calculated from simulations with an harmonic interaction, $\epsilon=10$, $\rho=0.05$, $T_{A}=1$ and different values of $T_{B}$ are presented in Fig. 1(a). As expected, we observe an increase of the correlation at contact $\hat{g}(1)$ due to the depletion interaction. This increase scales as $\epsilon^{2}$ with the interaction strength $\epsilon$, as expected from the different theoretical predictions (Fig. 1(b)). However, in contradiction with the predictions, the depletion interaction extends beyond two diameters as soon as the system departs from equilibrium. Beyond two particle diameters, the radial dependence of the depletion interaction is compatible with an algebraic decay, $\hat{g}(r)=G/r^{4}$. The prefactor $G$ of the algebraic decay increases with the interaction strength as $\epsilon^{3}$ for small $\epsilon$ and saturates for large values of $\epsilon$, corresponding to the hard sphere limit (Fig. 1(b)). The effect of the temperature $T_{B}$ is non-monotonic: $G$ first increases and then decreases at large temperatures (Fig. 1(c)). The temperature where the effect is maximal increases with the interaction strength $\epsilon$. We now return to theory to (i) elucidate the discrepancy between the two arguments adapted out of equilibrium and (ii) find the origin and the characteristics of the algebraic decay.
@@
$\hat{g}(r)=\frac{V*V(r)}{T_{A}T_{AB}}+\mathcal{O}(\epsilon^{3}).$ (9)
This expression is in quantitative agreement with the simulations for $r\lesssim 2$ (App. D.4.2, Fig. 5). It takes a similar form as the results obtained by adapting equilibrium arguments (Eqs. (6, 8)), but the temperatures enter differently in the prefactor. As the expression (9) is exact at the order $\epsilon^{2}$, we conclude that none of the two equilibrium calculations can be adapted out of equilibrium.
Figure 2: Prefactor $G$ of the algebraic decay as a function of the interaction strength $\epsilon$ for the harmonic (blue circles), Gaussian (green diamonds), and WCA (red squares) interactions from the numerical simulations. The lines represent the several asymptotic expressions: weak interaction for an integrable potential (black solid line), hard spheres (black dashed line), weak interaction for a WCA potential (red dash-dotted line) and strong interaction from a Gaussian interaction (green dash-dotted line).
Yet, the expression (9) vanishes at two particle diameters and does not explain the algebraic decay observed in the simulations (Fig. 1(a)). This is expected, because the simulations indicate that the algebraic decay arises at order $\epsilon^{3}$, while Eq. (9) is limited to order $\epsilon^{2}$. We now turn to a small density expansion that is valid for any interaction strength. The $N$-particle distribution $f(\bm{X})$ is stationary under the Smoluchowski equation describing the microscopic dynamics (1):
$\partial_{t}f=\sum_{i}\nabla_{i}\cdot\left[T_{i}\nabla_{i}f(\bm{X})+f(\bm{X})\sum_{j\neq i}\nabla V_{ij}(\bm{x}_{i}-\bm{x}_{j})\right].$ (10)
To isolate the many-body effects from the pair correlation obtained by Grosberg and Joanny [18], we write it as
@@
The depletion interaction encoded in $\hat{g}(r)$ can be computed from the three-body distribution $f(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3})$. Expanding $f$ in powers of $\epsilon$ and integrating over $\bm{x}_{3}$, we find (App. D.3.1)
$\hat{g}=\int\text{d}\bm{x}_{3}\left[v_{13}v_{23}-\phi^{(2)}-\frac{1}{2}\left(v_{13}^{2}v_{23}+v_{13}v_{23}^{2}\right)\right.\\ \left.+(v_{13}+v_{23})\phi^{(2)}-\phi^{(3)}\right]+\mathcal{O}(\epsilon^{4}).$ (13)
@@
To find $\phi^{(3)}$, we keep the terms of order $\epsilon^{3}$ in Eq. (12):
$\sum_{i}T_{i}\nabla_{i}^{2}\phi^{(3)}\\ =\sum_{\langle ij\rangle}\nabla v_{ij}\cdot\left[(T_{ij}+\tau_{ij})\nabla_{i}\phi^{(2)}-(T_{ij}-\tau_{ij})\nabla_{j}\phi^{(2)}\right],$ (15)
@@
which can be solved using the expression of $\phi^{(2)}$ obtained previously (App. D.2.2). Using $\phi^{(2)}$ and $\phi^{(3)}$ in the correction (13) leads the depletion interaction at order $\epsilon^{3}$. The resulting expression contains many terms involving double integrals in Fourier space, which can be performed numerically, yielding a good agreement with the $\hat{g}(r)$ obtained from simulations (App. D.3.3, Fig. 5). The small wavevector behavior in Fourier space leads to an algebraic decay:
$\hat{g}(r)\underset{r\to\infty}{\sim}G_{U}\left(\frac{\epsilon}{T_{A}},\frac{T_{B}}{T_{A}}\right)r^{-2d},$ (16)
where
$G_{U}(e,\theta)\underset{e\to 0}{\sim}\left[e\tilde{U}(0)\right]^{3}F(\theta).$ (17)
This prediction agrees quantitatively with the simulations for the harmonic and Gaussian interactions, which are integrable (Fig. 2). The temperature dependence follows $F(\theta)\sim\theta-1$ close to equilibrium ($\theta\to 1$), consistent with the fact that the algebraic interaction emerges from the temperature differences. When the $B$ particles are very hot ($\theta\to\infty$), they are barely affected by the interactions and the prefactor decays as $F(\theta)\sim\theta^{-2}$. These behaviors are confirmed by the simulations (Fig. 1(c)).
Furthermore, the simulations show that the algebraic decay as $r^{-2d}$ holds for the three different interactions, harmonic, WCA, Gaussian, independently of their strength (App. E.1, Fig. 6(a,b)). Hard spheres are obtained as the strong interaction limit for interactions with a finite range $\sigma$. In this limit, the prefactor depends only on the diameter $\sigma$ (Fig. 2). Assuming that the correlation depends only on the dimensionless ratio $r/\sigma$, the prefactor should thus be
$G_{U}(e,\theta)\underset{e\to\infty}{\to}\sigma^{2d}F_{\mathrm{hs}}(\theta);$ (18)
this argument is confirmed by simulations (Fig. 6(c)). Two situations remain uncovered by the analysis above: the WCA interaction in the soft limit, because it is not integrable, and the Gaussian interaction in the hard limit, because its range is not finite. They can be adressed by defining the effective diameter $\sigma_{\mathrm{eff}}$ as $\epsilon U(\sigma_{\mathrm{eff}})=T_{A}$, and then using it as an effective hard sphere diameter. For the WCA interaction in the limit $\epsilon\to 0$, this gives $\sigma_{\mathrm{eff}}^{\mathrm{WCA}}\sim(\epsilon/T_{A})^{1/12}$ and $G_{U}(e,\theta)\sim e^{1/3}$. For the Gaussian interaction in the limit $\epsilon\to\infty$, this gives $\sigma_{\mathrm{eff}}^{\mathrm{Gauss}}\sim\sqrt{\log(\epsilon/T_{A})}$ and $G_{U}(e,\theta)\sim[\log(e)]^{2}$. These behaviors are compatible with the simulations (Fig. 2). Finally, we ran numerical simulations with a Lennard-Jones interaction, which has an attractive part. Two behaviors are observed: if the attraction is small enough, the system remains homogeneous and the algebraic decay of the correlation is observed; if the attraction is too strong, the cold particles aggregate and the algebraic decay is lost (App. E.2, Fig. 7).
@@
The depletion interaction that we have unveiled is the driving force behind the formation of dense droplets of cold particles when they separate from the hot particles. As the algebraic tail of the interaction originates from temperature differences ($\tau_{ijk}$ in Eq. (12)), this effect is reversed if one considers the depletion interaction induced by the cold particles on the hot ones, so that the algebraic tail is repulsive in this case. This repulsion between hot particles may work with the attraction between cold particles to trigger the phase separation. Our results may also apply to mixtures of passive and artificial [38, 39] or living [40, 41] self-propelled particles, which could be described at large scales by two-temperature mixtures.